OpenStudy (anonymous):

Evaluate the definite integrals. \[\int\limits_{0}^{4} (\frac{ 3 }{ 2x+1 })dx\]

OpenStudy (anonymous):

u=2x+1 du=2dx ??

ganeshie8 (ganeshie8):

thats a good start

OpenStudy (anonymous):

New limit x=0, u=1 x=4, u=9

OpenStudy (anonymous):

So it would like this?.. \[\int\limits_{1}^{9}\frac{ 3 }{ u }du\] ??

ganeshie8 (ganeshie8):

not quite, there is a mistake with differential du=2dx dx = ?

OpenStudy (anonymous):

dx= 1/2 du ?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

Oh so it would be: \[\int\limits_{1}^{9}\frac{ 3 }{ 2u }du\]

OpenStudy (anonymous):

Then I got 3/2ln9 as the final answer..?

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

Is it possible to reduce that even more?

ganeshie8 (ganeshie8):

use the log property \[\large n\,\log x ~~=~~\log x^n \]

OpenStudy (anonymous):

I'm not sure how to use it correctly...

ganeshie8 (ganeshie8):

\[\large \frac{3}{2}\,\ln(9) ~~=~~\ln (9^{\frac{3}{2}}) \]

OpenStudy (anonymous):

Then what happens?...

ganeshie8 (ganeshie8):

\[\large \frac{3}{2}\,\ln(9) ~~=~~\ln (9^{\frac{3}{2}}) =\ln({\sqrt{9~}}^3)\]

OpenStudy (anonymous):

Oh I see! It would be 3ln3!

ganeshie8 (ganeshie8):

Yep!

OpenStudy (anonymous):

Thank you! :)

ganeshie8 (ganeshie8):

yw!